The probability that Kodjo and Adoga pass an examination are
3/4 and 3/5 respectively. Find the probability of both boys failing the examination.
1/3
3/10
9/20
1/3
1
Expert's answer
2015-02-25T13:00:57-0500
The probability of both boys failing the examination is (1-3/4)*(1-3/5)=1/4*2/5=1/10 by multiplication rule of probability.